OpenStudy (jangyj001):

Suppose we have the following information about how many Canadians were watching these events: 56% watched Ski Jumping (SK) 42% watched Snowboarding (SB) 67.44% watched Ski Jumping or Speed Skating (SS) 55.65% watched Snowboarding or Speed Skating 35.86% watched Ski Jumping and Snowboarding 14.56% watched Ski Jumping and Speed Skating 8.29% watched all three events. We will randomly select one Canadian. What is the probability that the selected person didn't watch any of the three events?

OpenStudy (kirbykirby):

Let: \(K\)=ski jumping \(B\)=snow boarding \(S\)=speed skating Then, we want to know the prob. that a person didn't watch any of the 3 events. This can be translated as "Didn't watch \(K\), and didn't watch \(B\), and didn't watch \(S\)". In probability terms, you want \(P(\overline{K} \cap \overline{B} \cap \overline{S})\) which equals to: \(P(\overline{K \cup B \cup S})\) by DeMorgan's Law. What we are given: \(P(K)=0.56\) \(P(B)=0.42\) \(P(K \cup S)=0.6744\) \(P(B \cup S)=0.5565\) \(P(K \cap B)=0.3586\) \(P(K \cap S)=0.1456\) \(P(K \cap B \cap S)=0.0829\) By the inclusion/exclusion principle: \[P(K \cup B \cup S)=P(K)+P(B)+\color{red}{P(S)}-P(K \cap B)\\~~~~~~~~~~~~~~~~~~~~~~~~~~~-P(K \cap S)-\color{red}{P(B \cap S)}+P(K \cap B \cap S)\] In red is the information not directly given, which you can find with other probability rules.

OpenStudy (kirbykirby):

\(P(B \cap S)\) can be found from: \(P(B \cup S)=P(B)+P(S)-P(B \cap S)\) Although you're still missing P(S), which could be found from: \(P(K \cup S)=P(K)+P(S)-P(K \cap S)\)